OpenStudy (isabel☺):

solve y lnx lny dx + dy =0

OpenStudy (dls):

\[\Large y lnx\ lny dx + dy =0\] \[\Large y lnx\ lny dx = -dy\] \[\Large lnx dx = \frac{-dy}{y \ln y}\] Integrate both sides \[\Large xlnx-x+c = \ln(\ln y)+C\] \[\Huge x \ln x =\ln (\ln y)+x+K\]

OpenStudy (dls):

oops it would be - ln (lny) forgot the negative sign :)

OpenStudy (isabel☺):

thanks i m so glad we have the same answer :D

OpenStudy (dls):

:)